View Zoom Add Page Insert Table Chart Text Shape Media Comment Collaborate Format Document Text Normal (Web) 023 4 X -1600 1200 300 300 Calculate NPV@14% & IRR 300 Style More 14. An investment is expected to pay you $5000 per year for $10 years. What is this investment worth if you have an 8% required return? Font 13. If interest rates rise, what will bond prices do? What type of bonds will show the biggest s New Roman change in price? (High oe low coupon? Long or short maturity?) 12 pt 16. If interest rates stay constant, how will a boads price change as time passes? 17. What is a callable bond? Character Styles None 18. A stock recently paid a $2 dividend. If that dividend is expected to grow by 10% per year, what should the stocks price be if you require a l % return? Alignment 19. A stock is expected to pay a $1.40 divided at the end of one year. If you expeet the dividend to 10%? grow by 6% per year, what should the stock price be if the required return is 20. A stock hasa beta of 120. If the risk-free rate is 2% and the market risk premium is 6% what does the CAPM say that the stocks required return should be? Spacing 10-Single 21. Standard deviation measures_Unsystematicisk. Beta measures Systemati risk. 22. Combining stocks with corrclation will create the most diversification. Bullets& Lists None (high/low) 23, You are considering an investment that will make 8% over the next year. However, inflation is expected to bc 3%. According to the Fisher Equation, what is your upproximate real returm on the investment? 24. You bought a stock last year for $55. During the year, it paid a $2 dividend. Youre selling it today for $62. What is your percent return on this investmem? 602 words
Answer #1

1. NPV = -1600 + 1200/(1+14%) + 300/(1+14%)2 + 300/(1+14%)​​​​​​3 +300/(1+14%)4 = 63.59
IRR using excel formula

A B C D E
1 -1600 1200 300 300 300
IRR 16.66% Excel formula = IRR(A1:E1)

2.
PV using annuity formula = PMT*(1-(1+r)-n)/r = 5000 *(1-(1+8%)-10)/8% = 33,550.41


3. IF interest rates rise Bond prices will fall.
Biggest change in price will be found in long maturity and low coupon bonds

4. If interest rates are same the bond prices  increases since maturity keeps on decreasing and hence par value is discounted for lower time.

Max 4 sub parts cane be solved.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well